From a point A, a body travels for 2s at a constant velocity of 12m/s due north. It then travels for 5s due east whilst its velocity decreases uniformly from 12m/s to 6m/s. What is the resultant displacement from A? Also determine the average velocity, in magnitude and direction, of the body for it to return along a straight path to point A in 3s.
1
Expert's answer
2019-04-12T09:39:02-0400
"v_{x,avg}=\frac{12+6}{2}=9 \frac{m}{s}"

"s_y=12(2)=24 m"

"s_x=9(5)=45 m"

"s=\sqrt{24^2+45^2}=51 m."

The direction is


"a=\arctan{\frac{24}{45}}=28^\circ"

north of east.

The direction, of the body for it to return along a straight path to point A in 3s is 28 degree south of west.

The magnitude is


"V=\frac{51}{3}=17 \frac{m}{s}"